Tiles with each of the 26 letters of the alphabet printed on them are placed in a box. If three tiles are selected from the box, with replacement, what is the probability that the tiles selected will spell the word ADD? the letters must be in order.
Answer by stanbon(57219)

Tiles with each of the 26 letters of the alphabet printed on them are placed in a box. If three tiles are selected from the box, with replacement, what is the probability that the tiles selected will spell the word ADD? the letters must be in order.
---
Since the drawing is made "with replacement" the probability
of drawing each letter is 1/26.
P(ADD) = (1/26)^3 = 1/17576
==================================
Cheers,
Stan H.